OpenStudy (anonymous):

solve for x. 6-(2-3 cot x)= 5 cot x

OpenStudy (anonymous):

6-2+3cot x =5cot x 4=2cot x cot x = 2 tan x =1/2 \[x =\tan^{-1} \frac{1}{2}\]

OpenStudy (anonymous):

oops, beat to the punch

OpenStudy (anonymous):

didn't we need to factor ?? or no?

OpenStudy (anonymous):

he factored the negative sign then added/subtracted

OpenStudy (anonymous):

why should we need a factor..its not a quadratic equarion

OpenStudy (anonymous):

probably thought the 6 needed to be factored into ( )

OpenStudy (anonymous):

or foiled rather

OpenStudy (anonymous):

ok I got it thanks !!!